My father is currently 13 times as old as my daughter. Next year, my father will be 11 times as old as my daughter. How many years old is my father now?
Thank you for your time. I hope you can get back to me soon.

Found 2 solutions by Theo, macston:
Answer by Theo(13342)

f is your father's age.
d is your daughter's age.
today, f = 13 * d
next year (f+1) = 11 * (d+1)
in the second equation, replace f with 13 * d to get:
next year 13 * d + 1 = 11 * (d + 1)
simplify to get:
13 * d + 1 = 11 * d + 11
subtract d from both sides of the equation and subtract 1 from both sides of the equation to get:
2 * d = 10
divide both sides of the equation by 2 to get:
d = 5
your daughter is 5 years old today and your father is 13 * 5 = 65 years old.
next year your daughter will be 6 and your father will be 66.
65 is equal to 13 * 5
66 is equal to 11 * 6
your solution is that your father is 65 years old today.



Answer by macston(5194)

F=father's age; D=daughter's age
F=13D
F+1=11(D+1) Substitute for F.
13D+1=11D+11 Subtract 11D+1 from each side
2D=10 Divide each side by 2.
D=5 Daughter is 5 years old.
F=13D=13(5)=65 ANSWER: Father is 65 years old.
CHECK:
F+1=11(D+1)
65+1=11(5+1)
66=11(6)
66=66